Please help me to solve this problem. A corn refining company produces corn gluten cattle feed at a variable cost of $76 per ton. If fixed costs are $110,000 per month and the feed sells for $126 per ton, how many tons must be sold for the company to have a monthly profit of $540,000?
Answer by mananth(16946)

Let the number of tons produced per month be x
126x- (110,000 + 76x)=540,000
126x-110,000-76x=540,000
50x=650,000
/50
x=650,000/50
x=13000 Tons